# Inference-Time Scaling of Discrete Diffusion Models via Importance Weighting and Optimal Proposal Design

### v3

@@ 5 Related Work
Discrete Diffusion Models. Discrete diffusion models (DDMs) were originally introduced in Austin et al. (2021); Sun et al. (2022); Campbell et al. (2022), grounded in the framework of continuous-time Markov chains (Norris, 1998). More recently, masked diffusion models (MDMs) (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024), a special case of DDMs, have shown strong performance in language modelling (Zhang et al., 2025; Nie et al., 2025). In addition, MDMs have achieved promising results in math reasoning (Zhao et al., 2025), image synthesis (Bai et al., 2024), code planning (Gat et al., 2024; Gong et al., 2025), and biological sequence generation (Campbell et al., 2024), yielding performance comparable to continuous diffusion (Rombach et al., 2022) and autoregressive models (Radford et al., 2019). In contrast to these approaches, which primarily study large-scale pretraining, our work focuses on test-time inference and post-training alignment (Uehara et al., 2025b), where access to training data is not available.
@@
Test-time Alignment of Discrete Diffusion Models. Existing alignment methods mainly fall into two categories: classifier guidance (Dhariwal & Nichol, 2021) and RL-based fine-tuning (Black et al., 2023). Although the score is ill-defined in discrete distributions, several works (Vignac et al., 2022; Nisonoff et al., 2024; Schiff et al., 2024) employ a first-order approximation to the target distribution, which resembles the insight underlying our approximated proposal $\text{Prop}_{\text{grad}}$ in Equation˜8. Beyond guidance approaches, sampling-based techniques have also demonstrated promising performance, including value-based sampling (Li et al., 2024), importance sampling (Guo et al., 2024), and iterative refinement strategies (Uehara et al., 2025a). While training-free and relatively efficient to deploy, these methods often face challenges in scalability and robustness. More recently, RL-based fine-tuning methods (Zekri & Boullé, 2025; Zhao et al., 2025; Gong et al., 2025) have gained significant traction, fueled by the remarkable success of Group Relative Policy Optimisation (GRPO) (Shao et al., 2024) in large language models (Guo et al., 2025). In parallel, steering-based (Rector-Brooks et al., 2024) approaches leveraging GFlowNets (Bengio et al., 2023) and direct backpropagation methods (Wang et al., 2024) have also demonstrated strong potential for test-time alignment. Distinct from these directions, our amortised proposal $\text{Prop}_{\text{amot}}$ introduces an alternative perspective for fine-tuning pretrained discrete diffusion models: it minimises the log-variance of importance weights, a criterion that has been rarely investigated in previous work.
@@
Sequential Monte Carlo for Generative Modelling. SMC has emerged as a versatile framework for probabilistic modelling, providing effective tools for sampling and inference across a wide range of applications, including particle filtering (Johansen, 2009), Bayesian experimental design (Ryan et al., 2016), and probabilistic planning (Piché et al., 2018). Most recently, SMC has been combined with diffusion models (Chen et al., 2024; He et al., 2025; Skreta et al., 2025; Wu et al., 2025), transforming it into a powerful neural sampler capable of drawing from complex Boltzmann distributions. These developments have also extended SMC’s reach to discrete domains, as demonstrated by Holderrieth et al. (2025); Lee et al. (2025). Beyond classical sampling tasks, SMC has further expanded to the improvement of generative models at test time. A seminal step in this direction was taken by Zhao et al. (2024), who introduced SMC as a principled probabilistic inference framework for addressing capability and safety challenges in large language models (LLMs). Subsequent works (Feng et al., 2024; Puri et al., 2025) successfully applied this idea to enhance mathematical reasoning in LLMs, while others explored its use in reward-guided adaptation of pretrained diffusion models (Wu et al., 2023a; Cardoso et al., 2023; Kim et al., 2025). Our work is most closely related to Singhal et al. (2025); Dang et al. (2025), who employ SMC for test-time alignment of discrete diffusion models. However, their approaches treat pretrained diffusion models as fixed proposal distributions. By contrast, we take a closer look at the role of proposal choice, systematically investigating its impact and providing empirical evidence for a key insight: proposals that better approximate the optimal, which minimises the variance of importance weights, consistently lead to better performance.
@@ Appendix A Abstract Proof and Derivations
A.1 Proof of Locally Optimal Proposal
@@
A.2 Proof of Log-Variance Minimisation Objective
@@ B.2 Continuous-Time Formulation of SMC
Remark. Proposition˜1 recovers the importance weights proposed in the SMC methods of Holderrieth et al. (2025); Ou et al. (2025). In those works, the intermediate target distribution is defined as a geometric interpolation between the base and target distributions: $\pi(x_{t})=p_{\mathrm{base}}^{t}(x_{t})p_{\mathrm{target}}^{1-t}(x_{t})$. The proposal rate matrix $R_{t}$ is then trained to satisfy the Kolmogorov forward equation. In contrast, we consider a different scenario: the pretrained model is available, but the intermediate target $\pi(x_{t})$ cannot be computed explicitly.
@@
By applying the forward Kolmogrov equation in Equation˜28, we have
@@ C.1 Sampling Schemes in Masked Diffusion Models
Notably, the ReMDM training objective is a reweighted version of the standard masked diffusion loss in Equation˜3. Thus, we can take a pretrained masked diffusion model, and use the ReMDM sampling in Equation˜41 for inference.
@@
C.2 Computing Importance Weight with Low-Confidence Sampling
@@
While this computation is straightforward for MDM and ReMDM sampling, it becomes tricky for low-confidence sampling. The difficulty arises because $p$ and $q$ rely on different denoisers, denoted as $\mu_{\theta}$ and $\mu_{\phi}$, respectively. If one strictly follows the rule in Equation˜42, the log-ratio often collapses to zero whenever
@@ D.1.2 Training Details of the Amortised Proposal
Synthetic Experiments. In this experiment, we take the MDLM (Sahoo et al., 2024) as the pretrained diffusion model. Finetuning is performed on a single NVIDIA A6000 GPU with a batch size of $32$ for MNIST and $128$ for MoG. The model is trained for $30$ epochs on MNIST and $20$ epochs on MoG, with $5$ optimisation steps per epoch. To avoid out-of-memory issues, we compute the loss over $10$ randomly sampled time steps $t$ instead of using gradient accumulation, and choose $M=10$ to estimate the reward. The Adam optimiser (Adam et al., 2014) is applied to train both the model and $F_{\psi}$, with a learning rate of $0.001$ for MoG and $0.0001$ for MNIST.
Language Modelling. This experiment closely follows Singhal et al. (2025). The pretrained language model used is MDLM [^1] https://huggingface.co/kuleshov-group/mdlm-owt (Sahoo et al., 2024), which is trained on the OpenWebText dataset. We perform full-parameter finetuning on a single NVIDIA A6000 GPU with a batch size of $32$. Training is conducted for $50$ epochs, with $5$ optimisation steps per epoch. To avoid the memory issue, at each optimisation step we compute the loss using one randomly selected time step $t$, together with a fixed $t=0$. During training, rewards are scaled by a factor of $20$, and estimated with $M=20$ Monte Carlo samples. Both the model and $F_{\psi}$ are optimised using Adam (Adam et al., 2014) with a learning rate of $0.0001$.
Biology Design. This experiment focuses on regulatory DNA sequence generation. We use the pretrained masked discrete diffusion model from Wang et al. (2024) which has been trained on a dataset of $\sim 700$k DNA sequences (Gosai et al., 2023). We perform full-parameter finetuning on a single NVIDIA RTX 3090 GPU. Training is conducted for 350 epochs, with 10 optimisation steps per epoch. We use a batch size of 64 and use a sampling mix of $0.9:0.1$ of on-policy from $q_{\phi}$ and off-policy samples from $p_{\theta}$. To manage memory usage, we do a gradient accumulation of the loss at each timestep before taking an optimisation step. Additionally, we only consider the final 50 of the 128 timesteps for loss calculation, following Wang et al. (2024). We also add a negative entropy term of the form $\sum_{x_{t-1}}q_{\phi}(x_{t-1}|x_{t})\log q_{\phi}(x_{t-1}|x_{t})$ to the loss with a coefficient of 2.52.5222We empirically observe that the best value for the entropy coefficient varies proportionally with the reward scaling factor maintaining a ratio of 0.0020.0030.002-0.003.. We observe empirically that it helps in preventing mode collapse during training. For rewards, we use a scaling factor of $1000$ and estimate it using just $M=1$ Monte Carlo sample. Both the model and $F_{\psi}$ are optimized using the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $1\times 10^{-5}$.
@@
During training, we adopt classifier-free guidance (Ho & Salimans, 2022) with a guidance scale of $5$, using the negative prompt “worst quality, low quality, low res, blurry, distortion, watermark, logo, signature, text, jpeg artifacts, sketch, duplicate, ugly, identifying mark”, following the inference script provided by Meissonic. Reward rescaling proves to be critical for stable optimisation (Liu et al., 2024). Specifically, we multiply the reward by a coefficient $\beta$, setting $\beta=100$ for both the Aesthetic Score and ImageReward, and $\beta=10,000$ for HPSv2. The coefficient is linearly annealed from 0 to its maximum value over the first $25$ epochs. For ImageReward and HPSv2, no KL regularisation between the fine-tuned and pretrained models is applied. However, for the Aesthetic Score, we observe that incorporating a KL term of the form $\mathbb{KL}(q_{\phi}(x_{t-1}|x_{t})||p_{\theta}(x_{t-1}|x_{t}))$ with a coefficient of $0.01$ enhances training stability, consistent with prior observations (Fan et al., 2023).

### v4

@@ 5 Related Work
Discrete Diffusion Models. Discrete diffusion models (DDMs) were originally introduced in Austin et al. (2021); Sun et al. (2022); Campbell et al. (2022), grounded in the framework of continuous-time Markov chains (Norris, 1998). More recently, masked diffusion models (MDMs) (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024), a special case of DDMs, have shown strong performance in language modelling (Zhang et al., 2025a; Nie et al., 2025). In addition, MDMs have achieved promising results in math reasoning (Zhao et al., 2025), image synthesis (Bai et al., 2024), code planning (Gat et al., 2024; Gong et al., 2025), and biological sequence generation (Campbell et al., 2024), yielding performance comparable to continuous diffusion (Rombach et al., 2022) and autoregressive models (Radford et al., 2019). In contrast to these approaches, which primarily study large-scale pretraining, our work focuses on test-time inference and post-training alignment (Uehara et al., 2025b), where access to training data is not available.
@@
Test-time Alignment of Discrete Diffusion Models. Existing alignment methods mainly fall into two categories: classifier guidance (Dhariwal and Nichol, 2021) and RL-based fine-tuning (Black et al., 2023). Although the score is ill-defined in discrete distributions, several works (Vignac et al., 2022; Nisonoff et al., 2024; Schiff et al., 2024) employ a first-order approximation to the target distribution, which resembles the insight underlying our approximated proposal $\text{Prop}_{\text{grad}}$ in Equation˜8. Alternatively, Rout et al. (2025) perform guidance on the embedding space, mitigating the issue of ill-defined gradients. Chen et al. (2025b) introduce reward-free guidance, analogous to classifier-free guidance (Ho and Salimans, 2022) but designed for masked diffusion models. Moreover, Tang et al. (2025) propose tree search guided finetuning, which is related to the searching-based scaling methods on continuous diffusion (Ma et al., 2025; Zhang et al., 2025b; Jain et al., 2025; Ramesh and Mardani, 2025). Beyond guidance approaches, sampling-based techniques have also demonstrated promising performance, including value-based sampling (Li et al., 2024), importance sampling (Guo et al., 2024), and iterative refinement strategies (Uehara et al., 2025a). While training-free and relatively efficient to deploy, these methods often face challenges in scalability and robustness. More recently, RL-based fine-tuning methods (Zekri and Boullé, 2025; Zhao et al., 2025; Gong et al., 2025) have gained significant traction, fueled by the remarkable success of Group Relative Policy Optimisation (GRPO) (Shao et al., 2024) in large language models (Guo et al., 2025). In parallel, steering-based (Rector-Brooks et al., 2024) approaches leveraging GFlowNets (Bengio et al., 2023) and direct backpropagation methods (Wang et al., 2024) have also demonstrated strong potential for test-time alignment. Distinct from these directions, our amortised proposal $\text{Prop}_{\text{amot}}$ introduces an alternative perspective for fine-tuning pretrained discrete diffusion models: it minimises the log-variance of importance weights, a criterion that has been rarely investigated in previous work.
@@
Sequential Monte Carlo for Generative Modelling. SMC has emerged as a versatile framework for probabilistic modelling, providing effective tools for sampling and inference across a wide range of applications, including particle filtering (Johansen, 2009), Bayesian experimental design (Ryan et al., 2016), and probabilistic planning (Piché et al., 2018). Most recently, SMC has been combined with diffusion models (Chen et al., 2024; He et al., 2025; Skreta et al., 2025; Wu et al., 2025; Ou et al., 2026), transforming it into a powerful neural sampler capable of drawing from complex Boltzmann distributions. These developments have also extended SMC’s reach to discrete domains, as demonstrated by Holderrieth et al. (2025); Lee et al. (2025). Beyond classical sampling tasks, SMC has further expanded to the improvement of generative models at test time. A seminal step in this direction was taken by Zhao et al. (2024), who introduced SMC as a principled probabilistic inference framework for addressing capability and safety challenges in large language models (LLMs). Subsequent works (Feng et al., 2024; Puri et al., 2025) successfully applied this idea to enhance mathematical reasoning in LLMs, while others explored its use in reward-guided adaptation of pretrained diffusion models (Trippe et al., 2022; Wu et al., 2023a; Cardoso et al., 2023; Dou and Song, 2024; Kim et al., 2025; Yoon et al., 2025; Chen et al., 2025a; Ren et al., 2025). Our work is most closely related to Singhal et al. (2025); Dang et al. (2025); Hasan et al. (2025), who employ SMC for test-time alignment of discrete diffusion models. However, their approaches treat pretrained diffusion models as fixed proposal distributions. By contrast, we take a closer look at the role of proposal choice, systematically investigating its impact and providing empirical evidence for a key insight: proposals that better approximate the optimal, which minimises the variance of importance weights, consistently lead to better performance.
@@ 6 Conclusion
7 Acknowledgements
ZO is supported by the Lee Family Scholarship. We would like to thank the anonymous reviewers for their constructive and valuable suggestions. ZO also thanks Ruixiang Zhang for insightful discussions and for his support with the experimental setup.
@@ Appendix A Abstract Proof and Derivations
A.1 A Brief Recap of SMC
In this section, we provide a brief overview of Sequential Monte Carlo (SMC). For a target distribution $\pi(x_{t})$, we consider the problem of estimating the expectation of a test function $\delta$, namely ${\mathbb{E}}_{\pi(x_{t})}[\delta(x_{t})]$. When $\delta(\cdot)$ is taken to be the Dirac delta function, estimating this expectation reduces to constructing an empirical approximation of the distribution $\pi(x_{t})$.
To estimate the expectation, importance sampling introduces a proposal distribution $q$, which is easy to sample from, and proposes an estimator as follows
${\mathbb{E}}_{\pi(x_{t})}[\delta(x_{t})]\!=\!{\mathbb{E}}_{q(x_{t:T})}\left[\frac{\pi(x_{t:T})}{q(x_{t:T})}\delta(x_{t})\right]\!\approx\!\sum_{i=1}^{N}w_{t}^{(i)}\delta(x_{t}^{(i)}),\ \text{where}\ w_{t}^{(i)}\!=\!\frac{\pi(x_{t:T}^{(i)})}{q(x_{t:T}^{(i)})},x_{t:T}^{(i)}\sim q(x_{t:T})$
The key ingredients of SMC are the target distribution $\pi(x_{t:T})$ and the proposal distribution $q(x_{t:T})$. Here we consider the target distribution as a Markovian model associated with a sequence of forward transition kernels $\gamma$: $\pi(x_{t:T})=\pi(x_{t})\prod_{s=t}^{T-1}\gamma(x_{t+1}|x_{t})$; and the proposal distribution as $q(x_{t:T})=\pi(x_{T})\prod_{s=t}^{T-1}q(x_{t}|x_{t+1})$. Substituting these into the importance weights gives
$w_{t}$ $=\frac{\pi(x_{t})\prod_{s=t}^{T-1}\gamma(x_{t+1}|x_{t})}{\pi(x_{T})\prod_{s=t}^{T-1}q(x_{t}|x_{t+1})}$
$=\frac{\pi(x_{t})\gamma(x_{t+1}|x_{t})}{\pi(x_{t+1})q(x_{t}|x_{t+1})}\frac{\pi(x_{t+1})\prod_{s=t+1}^{T-1}\gamma(x_{t+1}|x_{t})}{\pi(x_{T})\prod_{s=t+1}^{T-1}q(x_{t}|x_{t+1})}$
$=\frac{\pi(x_{t})\gamma(x_{t+1}|x_{t})}{\pi(x_{t+1})q(x_{t}|x_{t+1})}w_{t+1}$ (13)
This recursive structure allows importance weights to be computed incrementally. SMC augments this with a resampling step to mitigate weight degeneracy. For $N$ particles, SMC proceeds as follows:
  • 1.
    Initialise: $x_{T}^{(i)}\sim\pi(x_{T}),w_{T}^{(i)}=1$.
  • 2.
    For $t=T,\dots,1$:
    • (a)
      Propagate: $x_{t-1}^{(i)}\sim q(x_{t-1}|x_{t}^{(i)})$.
    • (b)
      Update weights: $\frac{\pi(x_{t}^{(i)})\gamma(x_{t+1}^{(i)}|x_{t}^{(i)})}{\pi(x_{t+1})q(x_{t}^{(i)}|x_{t+1}^{(i)})}$.
    • (c)
      Resample particles according to $\left\{\frac{w_{t-1}^{(i)}}{\sum_{j=1}^{N}w_{t-1}^{(j)}}\right\}_{i=1}^{N}$; then reset all weights to $w_{t-1}^{(i)}=1$.
The resulting set of particles $\{x_{0}^{(i)},w_{0}^{(i)}\}_{i=1}^{N}$ forms an empirical approximation of the target $\pi(x_{0})$.
A.2 Proof of Locally Optimal Proposal
@@
A.3 Proof of Log-Variance Minimisation Objective
@@ B.2 Continuous-Time Formulation of SMC
Remark. Proposition˜1 recovers the importance weights proposed in the SMC methods of Holderrieth et al. (2025); Ou et al. (2025). In those works, the intermediate target distribution is defined as a geometric interpolation between the base and target distributions: $\pi(x_{t})\propto p_{\mathrm{base}}^{t}(x_{t})p_{\mathrm{target}}^{1-t}(x_{t})$. The proposal rate matrix $R_{t}$ is then trained to satisfy the Kolmogorov forward equation. In contrast, we consider a different scenario: the pretrained model is available, but the intermediate target $\pi(x_{t})$ cannot be computed explicitly. Moreover, the importance weight in Equation˜32 can also be derived via the Radon–Nikodym derivative (Campbell et al., 2024, Appendix C.1), (Denker et al., 2025, Lemma 4). We instead adopt a discrete-time formulation and present a streamlined derivation to keep the exposition accessible for readers who may not be familiar with Radon–Nikodym derivative or path-measure theory.
@@
By applying the forward Kolmogrov equation in Equation˜29, we have
@@ C.1 Sampling Schemes in Masked Diffusion Models
Notably, the ReMDM training objective is a reweighted version of the standard masked diffusion loss in Equation˜3. Thus, we can take a pretrained masked diffusion model, and use the ReMDM sampling in Equation˜42 for inference.
@@
C.2 First-Order Approximation with Gumbel-Softmax Relaxation
To apply the first-order approximately optimal proposal, we need to compute $\nabla_{x_{t}}\hat{r}(x_{t})$, where $\hat{r}(x_{t})=\frac{1}{M}\sum_{m=1}^{M}r(x_{0}^{(m)}),\quad x_{0}^{(m)}\sim p_{\theta}(x_{0}|x_{t})$ as defined in Equation˜12. However, because $x_{0}^{(m)}$ is drawn via categorical sampling, $\hat{r}(x_{t})$ is not differentiable with respect to $x_{t}$. To address this, we use the Gumbel–Softmax reparameterization trick to obtain a differentiable surrogate. Concretely, we break the computation of $\hat{r}(x_{t})$ into three steps:
  • 1.
    compute the denoising logits: $p=\mu_{\theta}(x_{t})$, where $\mu_{\theta}$ is the denoising model;
  • 2.
    sample $x_{0}$: $x_{0}^{(m)}\sim\mathrm{Cat}(x;p)$;
  • 3.
    evaluate the reward: $\hat{r}(x_{t})=\frac{1}{M}\sum_{m=1}^{M}r(x_{0}^{(m)})$.
Following Grathwohl et al. (2021), we treat both $r$ and $\mu_{\theta}$ as functions that accept continuous inputs so that their gradients are well defined (steps 1 and 3). For step 2, we replace the categorical draw with its Gumbel–Softmax relaxation (Jang et al., 2016), making the sample $x_{0}^{(m)}$ differentiable with respect to $p$. Using these relaxations, the gradient can be obtained by the chain rule:
$\nabla_{x_{t}}\hat{r}(x_{t})\approx\frac{1}{M}\sum_{m=1}^{M}\frac{\partial r(x_{0}^{(m)})}{\partial x_{0}^{(m)}}\frac{\partial x_{0}^{(m)}}{\partial p}\frac{\partial p}{\partial x_{t}},\quad x_{0}^{(m)}\sim p_{\theta}(x_{0}|x_{t}).$ (44)
The first and last factors are provided by the differentiability of $r$ and $\mu_{\theta}$, while the middle term is approximated using the Gumbel–Softmax relaxation.
Remark. The gradient approximation in Equation˜44 requires the reward $r(x_{0})$ to be differentiable w.r.t. $x_{0}$. In the setting of non-differentiable rewards, one can apply the REINFORNCE gradient estimator (Williams, 1992). Specifically, we have
$\nabla_{x_{t}}\hat{r}(x_{t})$ $=\nabla_{x_{t}}{\mathbb{E}}_{p_{\theta}(x_{0}|x_{t})}[r(x_{0})]$
$=\sum_{x_{0}}p_{\theta}(x_{0}|x_{t})\nabla_{x_{t}}\log p_{\theta}(x_{0}|x_{t})r(x_{0})$
$={\mathbb{E}}_{p_{\theta}(x_{0}|x_{t})}[\nabla_{x_{t}}\log p_{\theta}(x_{0}|x_{t})r(x_{0})].$
Thus, the gradient can be approximated via
$\nabla_{x_{t}}\hat{r}(x_{t})\approx\frac{1}{M}\sum_{m=1}^{M}r(x_{0}^{(m)})\nabla_{x_{t}}\log p_{\theta}(x_{0}^{(m)}|x_{t}),\quad x_{0}^{(m)}\sim p_{\theta}(x_{0}|x_{t}).$ (45)
C.3 Computing Importance Weight with Low-Confidence Sampling
@@
While this computation is straightforward for MDM and ReMDM sampling, it becomes tricky for low-confidence sampling. The difficulty arises because $p$ and $q$ rely on different denoisers, denoted as $\mu_{\theta}$ and $\mu_{\phi}$, respectively. If one strictly follows the rule in Equation˜43, the log-ratio often collapses to zero whenever
@@ D.1.2 Training Details of the Amortised Proposal
Synthetic Experiments. In this experiment, we take the MDLM (Sahoo et al., 2024) as the pretrained diffusion model. Finetuning is performed on a single NVIDIA A6000 GPU with a batch size of $32$ for MNIST and $128$ for MoG. The model is trained for $30$ epochs on MNIST and $20$ epochs on MoG, with $5$ optimisation steps per epoch. To avoid out-of-memory issues, we compute the loss over $10$ randomly sampled time steps $t$ instead of using gradient accumulation, and choose $M=10$ to estimate the reward. The Adam optimiser (Adam and others, 2014) is applied to train both the model and $F_{\psi}$, with a learning rate of $0.001$ for MoG and $0.0001$ for MNIST.
Language Modelling. This experiment closely follows Singhal et al. (2025). The pretrained language model used is MDLM [^1] https://huggingface.co/kuleshov-group/mdlm-owt (Sahoo et al., 2024), which is trained on the OpenWebText dataset. We perform full-parameter finetuning on a single NVIDIA A6000 GPU with a batch size of $32$. Training is conducted for $50$ epochs, with $5$ optimisation steps per epoch. To avoid the memory issue, at each optimisation step we compute the loss using one randomly selected time step $t$, together with a fixed $t=0$. During training, rewards are scaled by a factor of $20$, and estimated with $M=20$ Monte Carlo samples. Both the model and $F_{\psi}$ are optimised using Adam (Adam and others, 2014) with a learning rate of $0.0001$.
Biology Design. This experiment focuses on regulatory DNA sequence generation. We use the pretrained masked discrete diffusion model from Wang et al. (2024) which has been trained on a dataset of $\sim 700$k DNA sequences (Gosai et al., 2023). We perform full-parameter finetuning on a single NVIDIA RTX 3090 GPU. Training is conducted for 350 epochs, with 10 optimisation steps per epoch. We use a batch size of 64 and use a sampling mix of $0.9:0.1$ of on-policy from $q_{\phi}$ and off-policy samples from $p_{\theta}$. To manage memory usage, we do a gradient accumulation of the loss at each timestep before taking an optimisation step. Additionally, we only consider the final 50 of the 128 timesteps for loss calculation, following Wang et al. (2024). We also add a negative entropy term of the form $\sum_{x_{t-1}}q_{\phi}(x_{t-1}|x_{t})\log q_{\phi}(x_{t-1}|x_{t})$ to the loss with a coefficient of 2.52.5222We empirically observe that the best value for the entropy coefficient varies proportionally with the reward scaling factor maintaining a ratio of 0.0020.0030.002-0.003.. We observe empirically that it helps in preventing mode collapse during training. For rewards, we use a scaling factor of $1000$ and estimate it using just $M=1$ Monte Carlo sample. Both the model and $F_{\psi}$ are optimized using the AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of $1\times 10^{-5}$.
@@
During training, we adopt classifier-free guidance (Ho and Salimans, 2022) with a guidance scale of $5$, using the negative prompt “worst quality, low quality, low res, blurry, distortion, watermark, logo, signature, text, jpeg artifacts, sketch, duplicate, ugly, identifying mark”, following the inference script provided by Meissonic. Reward rescaling proves to be critical for stable optimisation (Liu et al., 2024). Specifically, we multiply the reward by a coefficient $\beta$, setting $\beta=100$ for both the Aesthetic Score and ImageReward, and $\beta=10,000$ for HPSv2. The coefficient is linearly annealed from 0 to its maximum value over the first $25$ epochs. For ImageReward and HPSv2, no KL regularisation between the fine-tuned and pretrained models is applied. However, for the Aesthetic Score, we observe that incorporating a KL term of the form $\mathbb{KL}(q_{\phi}(x_{t-1}|x_{t})||p_{\theta}(x_{t-1}|x_{t}))$ with a coefficient of $0.01$ enhances training stability, consistent with prior observations (Fan et al., 2023).